OpenStudy (anonymous):

Solve using Elimination. 7x + 8y =19 3x +9y = (-3)

OpenStudy (anonymous):

are you solving for x or y?

OpenStudy (anonymous):

Both.

OpenStudy (anonymous):

x=-8y/7+19/7 y=-7x/8+19/8

OpenStudy (anonymous):

that might be wrong bcuz i got all confused but thats what i got

OpenStudy (anonymous):

Okay. Thanx

OpenStudy (displayerror):

You can solve for x or y first--either way, you get the same answer at the end. Using the first equation, we can solve for y in terms of x: \[7x + 8y = 19\] \[8y = 19-7x\] \[y = \frac{19-7x}{8}\] Now that we have y in terms of x, we can plug into the second equation to eliminate y and solve for a value of x: \[3x + 9y = -3\] \[3x + 9 \left( \frac{19-7x}{8} \right) = -3\] Multiply by 8 on both sides or find the LCD to combine the fractions. I'll multiply by 8 here. \[8 \times \left( 3x + 9 \left( \frac{19-7x}{8} \right) \right) = 8 \times -3\] When we distribute, we can remove the 8 from the denominator so that we don't have to deal with fractions. Remember to multiply both sides of the equation, though! \[24x + 9 \left( 19-7x \right) = -24\] Distribute the 9 \[24x + 171 - 63x = -24\] You should now be able to solve for x by combining like terms. Now that we know what x is, we can plug back into the expression we solved for in the beginning that gives us what y is in terms of x: \[y = \frac{19-7x}{8}\] Then we will have solved for x and y.

OpenStudy (anonymous):

Thnx for the explanation too